# Protap: A Benchmark for Protein Modeling on Realistic Downstream Applications

### v2

@@ A.1 Descriptions and Categorization of Domain-Specific Models
The primary distinction between domain models and general models lies in their task-specific design. Domain models are tailored for particular tasks by enabling interaction between representations of different components and incorporating biochemical priors to enhance the encoder. In addition, domain models also leverage knowledge from other biochemical databases to preprocess the data, enabling the extraction of task-relevant key regions. Below, we summarize the characteristics of each domain model in relation to its respective tasks. Detailed task descriptions and model architectures are found in Appendix C and Appendix D.
Table 9: Domain Model Categorization
@@
  • UniZyme. UniZyme [22] not only incorporates pretraining for enzyme active site prediction but also introduces energetic frustration, an intrinsic biophysical phenomenon that directly reflects active site properties and catalytic mechanisms. The integration of such biochemical priors enhances the generalization capability of the enzyme encoder.
  • ClipZyme. ClipZyme [13] formulates enzyme function prediction as a reaction-centric retrieval task, aligning enzyme representations with chemical reaction embeddings in a shared latent space. By explicitly modeling atom-mapped reaction graphs and constructing pseudo-transition states, the framework integrates reaction mechanism information that is specific to enzymatic catalysis.
@@
  • DeepPROTACs. DeepPROTACs [23] circumvent explicit modeling of the ternary complex by encoding different components of the Target protein–PROTAC–E3 ligase system using separate neural network modules.
  • ET-PROTACs. ET-PROTACs [24] utilizes a cross-modal strategy and ternary attention mechanism, the model fully accounts for the cross-talk between PROTACs, target proteins, and E3 ligases, enabling more accurate modeling of ternary complex interactions.
@@
  • KDBNet. KDBNet [43] does not model the entire protein, instead, it defines the binding pocket based on prior knowledge from the KLIFS database. In modeling kinases, it incorporates both geometric and evolutionary features, including backbone torsion angles and embeddings derived from the ESM language model.
  • MONN. MONN [44] is a multi-objective neural network designed to simultaneously predict non-covalent interactions and binding affinity between compounds and proteins, without relying on structural information.
@@
  • DeepFRI. DeepFRI [45] employs graph convolutional networks (GCNs) [59] to process protein contact maps and integrates representations pretrained on Pfam sequences.
  • DPFunc. DPFunc [21] identifies key functional regions within protein structures and precisely predicts their associated biological functions by leveraging protein domain information.
A.2 Additional Observations
Table 10: Results on KIBA Dataset.
@@ A.2 Additional Observations
EGNN 0.434 $\pm$ 0.007 65.09 $\pm$ 0.68
w/MLM 0.398 $\pm$ 0.007 65.75 $\pm$ 0.65
w/MVCL 0.401 $\pm$ 0.003 64.69 $\pm$ 0.17
w/PFP 0.396 $\pm$ 0.006 65.27 $\pm$ 0.15
SE(3) Trans 0.442 $\pm$ 0.034 63.19 $\pm$ 1.15
w/MLM 0.619 $\pm$ 0.017 42.36 $\pm$ 2.52
w/MVCL 0.473 $\pm$ 0.026 58.57 $\pm$ 1.84
w/PFP 0.675 $\pm$ 0.007 40.99 $\pm$ 1.07
ProtBERT 0.471 $\pm$ 0.005 60.89 $\pm$ 3.66
w/MLM 0.409 $\pm$ 0.004 63.72 $\pm$ 0.83
w/MVCL 0.393 $\pm$ 0.001 65.10 $\pm$ 0.76
w/PFP 0.419 $\pm$ 0.003 64.45 $\pm$ 0.11
ESM 0.454 $\pm$ 0.004 58.14 $\pm$ 0.01
Table 11: Performance of SaProt across different applications.
PCS (C14.005) PCS (M10.003) PROTACs MTP
Model AUC (%)↑ AUPR (%)↑ AUC (%)↑ AUPR (%)↑ Acc (%)↑ AUC (%)↑ Pear ↑ AUC (%)↑
SaProt 96.16 $\pm$ 0.16 42.91 $\pm$ 0.27 89.18 $\pm$ 0.06 13.48 $\pm$ 0.03 73.44 $\pm$ 1.32 80.29 $\pm$ 0.70 0.492 76.89
We additionally report the performance of EGNN, SE(3)-Transformer, ProteinBERT, and ESM on the PLI task using the KIBA dataset, as shown in Tab. 10. Compared with the result on the DAVIS dataset, we observe that as the dataset size increases from 14,464 to 89,958, the performance of all models on the PLI task improves to varying degrees. Notably, the SE(3) Transformer benefits little from pretraining, whereas ProteinBERT shows a moderate performance gain due to pretraining.
Observation 1. Model performance on downstream tasks like PLI consistently improves with increased data size, indicating a power-law-like scaling trend.
Moreover, consistent with the fine-tuning strategy used for ESM2, we freeze the parameters of SaProt and leverage its representations to fine-tune and evaluate on two downstream tasks: PCS and PROTACs. We also report its performance on the mutation effect prediction task.
Observation 2. After integrating the Structure-Vocabulary, SaProt achieves strong performance on most downstream tasks. However, its results on the PCS and PROTACs benchmarks remain slightly below those of D-Transformer and EGNN, indicating that PCS and PROTACs continue to be challenging evaluation standards for protein language models. In contrast, SaProt greatly outperforms other models in mutation effects prediction, including ESM, with a comparable parameter. This gain is likely attributable to the Structure-Vocabulary, which reduces local sequence ambiguity and mitigates semantic confusion.
Furthermore, we visualize the relationship between model size and performance across four downstream tasks, as shown in Fig. 2 and Fig. 3. Our observation is summarized as follows:
Observation 3. Large-scale models such as SaProt and ESM2 demonstrate clear advantages on tasks that require only a single protein sequence as input, such as PFA and MTP. However, when downstream tasks involve more complex inputs, e.g., the PROTACs task, which implicitly requires the alignment between protein and compound representations, smaller models like D-Transformer and EGNN tend to perform better, suggesting their stronger capacity to handle multi-modal integration despite having fewer parameters.
Observation 4. For domain-specific models, we do not observe clear evidence that larger model size consistently leads to better performance across downstream tasks. In the PROTACs task, both domain-specific models perform poorly. However, in the PLI and GO tasks, these models achieve strong results despite their relatively small sizes, suggesting that domain knowledge may play a more critical role than scale in certain contexts.
@@
Figure 2: Visualization of model size and performance across two specialized tasks: (a) PCS and (b) PROTACs. Each point represents a model, with size indicating parameter count and color intensity reflecting task performance. Each point represents a model, with its size corresponding to the number of parameters, ranging from 0.2M for GVP to 650M for ESM. Larger points indicate larger models. The color intensity of each point indicates performance, with darker shades representing stronger results. The dashed circular boundaries group models by category: models incorporating geometric information in color, protein language models in color, and domain-specific models in color.
@@
Figure 3: Visualization of model size and performance across two general tasks: (a) PLI and (b) PFA. Each point represents a model, with size indicating parameter count and color intensity reflecting task performance.
Appendix B Preliminaries
B.1 Protein Data
@@
Sequence. The most common representation is the amino acid sequence, typically in FASTA format using single-letter codes. Sequence-based models, originally RNNs and CNNs, are now dominated by Transformers due to their superior long-range modeling. Large pre-trained models such as ESM [16, 17], ProteinBERT [18], and ProtTrans [60] capture biochemical properties and evolutionary relationships, with structural information implicitly encoded in their learned representations.
2D Graph.. Proteins can also be described as molecular graphs, with residues as nodes and chemical or spatial relationships as edges. 2D graphs capture connectivity and topological features but lack explicit spatial details. Contact maps and models like GraphFormer [14] encode residue-residue proximity or graph connectivity into the attention mechanism, partially reflecting structural information.
3D Structure. Many methods have been developed for modeling the 3D structures of proteins. 3D CNNs employ convolutional neural networks to extract local spatial features of proteins. Graph neural networks are widely used for representing protein and molecular structures, with examples including proteinMPNN [12], SE(3)-Transformer [41], GVP [10], and GearNet [19]. GearNet applies contrastive learning for the pre-training of protein structures. Additionally, Transformer-M [15] incorporates 3D information through distance matrices, enabling transformers to represent protein 3D structures while maintaining invariance. More recently, some works have begun to include protein surface information in modeling; for instance, ProteinINR [61] integrates sequence, 3D structure, and surface information to enhance protein representations.
B.2 Terminology
@@
Protein-Ligand Binding Affinity. Protein–ligand binding denotes the selective, high-affinity association between a protein and a small-molecule ligand to form a stable complex. Binding strength is quantified by the change in Gibbs free energy ($\Delta$G) or related measures such as the dissociation constant (Kd); a more negative $\Delta$G (or lower Kd) reflects stronger affinity. Reliable computational estimation of binding affinity accelerates virtual screening and lead optimization in drug discovery by prioritizing ligands most likely to bind their target proteins.
@@
Appendix C Detail of Downstream Application and Dataset
C.1 Enzyme-Catalyzed Protein Cleavage Site Prediction
Task Definition. Cleavage site prediction is formulated as a large-scale, imbalanced, residue-level binary classification problem. Let $P_{s}$ denote a substrate protein of length $\lvert R\rvert$, and for each residue $r_{j}$ ($j=1,\dots,\lvert R\rvert$), let $x_{j}\in\mathbb{R}^{d}$ be its feature vector. The objective is to learn a function
@@ C.1 Enzyme-Catalyzed Protein Cleavage Site Prediction
that outputs $1$ if $r_{j}$ is a cleavage site under catalysis by a given enzyme, and 00 otherwise. Formally, each substrate $P_{s}$ is associated with a label vector $Y\in\{0,1\}^{\lvert R\rvert},$ where $Y_{j}=1$ indicates cleavage at residue $j$. Training proceeds over residue-level examples $\{(x_{j},\,Y_{j})\}$ sampled from annotated cleavage datasets, enabling model generalization to varied substrate contexts.
@@
  • Data Source. The original data are from the MEROPS [57] database. This dataset has a public website, available at the following address: http://cadd.zju.edu.cn/protacdb/
@@
    • Prior studies [57] have shown that slight sequence variations among enzymes within the same MEROPS category are negligible. As a result, we generalize the hydrolysis data from a specific substrate-enzyme interaction to all enzymes in that category. This allowed us to augment our dataset by linking each substrate not only to its originally annotated enzyme but also to other enzymes classified in the same MEROPS group. Finally, we selected two enzyme families, C14.003 and M10.003, for evaluation.
@@
    Table 12: Statistics of the dataset after preprocessing.
    Enzyme Family #Train #Test Cleavage Sites (Train) Cleavage Sites (Test)
@@
  • Hosting. A preprocessed version of the dataset can be downloaded from https://huggingface.co/datasets/findshuo/Protap/tree/main/PCSP.
@@
C.2 Targeted Protein Degradation by Proteolysis-Targeting Chimeras
@@
Settings. For this task, we retain the same architectural hyperparameter configuration as used during pretraining. To encode SMILES representations, we adopt the GVP [10] architecture with a node dimensionality of 128, an edge dimensionality of 32, and a total of 3 layers. Models are trained for 50 epochs with a learning rate of 5e-4, a batch size of 24, and the Adam optimizer. We report the mean and standard deviation of performance across a fixed set of random seeds. All experiments are conducted on NVIDIA L40 GPUs, with per-run training time ranging from approximately 20 minutes to 6 hours, depending on the model and task complexity.
@@
  • Data Source. The original data was collected from the PROTAC-DB [58]. This dataset has a public website, available at the following address: http://cadd.zju.edu.cn/protacdb/
@@
    • Following the processing strategy proposed by ET-PROTACs [24], we assign a degradation label to each triplet in the dataset. Specifically, we first examine whether a PROTAC entry contains DC50 measurements. If available, entries with DC50 values below 1000 nM are labeled as active, while those equal to or above 1000 nM are labeled as inactive. For entries lacking DC50 data, we instead inspect the reported degradation percentage of the target protein. If available, samples with degradation rates of at least 70% are considered active, and those below 70% are labeled as inactive. In cases where neither DC50 nor degradation percentage is reported, we fall back to IC50 values. Similarly, entries with IC50 below 1000 nM are labeled as active, and those above or equal to this threshold are labeled as inactive.
    • For the target protein and E3 ligase, we retained only the entries with available structures in the AFDB. We filtered out the target proteins with the following UniProt IDs: P36969, P03436, P0DTD1, Q12830, and G0R7E2. The structural data of the target proteins and E3 ligases are obtained from AFDB. Following the approach in [62], we further convert the format into a dictionary and store it in a .json file. The fields are as follows:
@@
                  "seq": "MADEEAGGTERMEISAELPQTPQRLASWWDQQVDFYTA...",
                  "coord": [[[21.9960, 68.3170,-49.9029], ...],...]
[…]
@@
    uniprot  e3_ligase_structure linker_sdf     warhead_sdf     e3_ligand_sdf    label
    Q00987   Q96SW2              linker_2.sdf   warhead_7.sdf   e3_ligand_7.sdf    1
    Q00987   Q96SW2              linker_2.sdf   warhead_7.sdf   e3_ligand_16.sdf   1
    P10275   P40337              linker_13.sdf  warhead_27.sdf  e3_ligand_27.sdf   0
    P10275   P40337              linker_33.sdf  warhead_27.sdf  e3_ligand_27.sdf   0
@@
    Table 13: Statistics of PROTACs dataset components.
    Object Target Protein E3-Ligase Warhead Linker E3-Ligand Label: 1 Label: 0
    Number 154 10 552 1,534 131 2,244 1,878
  • Hosting. A preprocessed version of the dataset can be downloaded from https://huggingface.co/datasets/findshuo/Protap/tree/main/PROTACs.
@@
C.3 Protein–Ligand Interactions

### v3

@@ D.1. Descriptions and Categorization of Domain-Specific Models
The primary distinction between domain models and general models lies in their task-specific design. Domain models are tailored for particular tasks by enabling interaction between representations of different components and incorporating biochemical priors to enhance the encoder. In addition, domain models also leverage knowledge from other biochemical databases to preprocess the data, enabling the extraction of task-relevant key regions. Below, we summarize the characteristics of each domain model in relation to its respective tasks. Detailed task descriptions and model architectures are found in Appendix F and Appendix H.
Table 8. Domain Model Categorization
@@
  • UniZyme. UniZyme (Li et al., 2025) not only incorporates pretraining for enzyme active site prediction but also introduces energetic frustration, an intrinsic biophysical phenomenon that directly reflects active site properties and catalytic mechanisms. The integration of such biochemical priors enhances the generalization capability of the enzyme encoder.
  • ClipZyme. ClipZyme (Mikhael et al., 2024) formulates enzyme function prediction as a reaction-centric retrieval task, aligning enzyme representations with chemical reaction embeddings in a shared latent space. By explicitly modeling atom-mapped reaction graphs and constructing pseudo-transition states, the framework integrates reaction mechanism information that is specific to enzymatic catalysis.
@@
  • DeepPROTACs. DeepPROTACs (Li et al., 2022) circumvent explicit modeling of the ternary complex by encoding different components of the Target protein–PROTAC–E3 ligase system using separate neural network modules.
  • ET-PROTACs. ET-PROTACs (Cai et al., 2025) utilizes a cross-modal strategy and ternary attention mechanism, the model fully accounts for the cross-talk between PROTACs, target proteins, and E3 ligases, enabling more accurate modeling of ternary complex interactions.
@@
  • KDBNet. KDBNet (Luo et al., 2023b) does not model the entire protein, instead, it defines the binding pocket based on prior knowledge from the KLIFS database. In modeling kinases, it incorporates both geometric and evolutionary features, including backbone torsion angles and embeddings derived from the ESM language model.
  • MONN. MONN (Li et al., 2020b) is a multi-objective neural network designed to simultaneously predict non-covalent interactions and binding affinity between compounds and proteins, without relying on structural information.
@@
  • DeepFRI. DeepFRI (Gligorijević et al., 2021) employs graph convolutional networks (GCNs) (Kipf and Welling, 2017) to process protein contact maps and integrates representations pretrained on Pfam sequences.
  • DPFunc. DPFunc (Wang et al., 2025) identifies key functional regions within protein structures and precisely predicts their associated biological functions by leveraging protein domain information.
D.2. Detail of Friedman–Nemenyi Tests for Pretraining Objectives
We compare three pretraining strategies (MLM, MVCL, PFP) using rank-based Friedman tests and Nemenyi post-hoc analyses across five backbones (EGNN, SE(3)-Trans, GVP, ProtBERT, D-Trans). To avoid redundancy across correlated metrics, we select one representative metric per task: AUC for PCS (C14.005) and PROTACs, Pearson correlation for PLI, and $F_{\max}$ for PFA, yielding $5\times 4=20$ evaluation blocks.(MTP is excluded since it is only applicable to MLM). Lower average rank is better.
Table 9. Statistical comparison of pretraining objectives.
@@ D.2. Detail of Friedman–Nemenyi Tests for Pretraining Objectives
Overall Mixed 20 6.10 0.047 0.153 0.741 1.75 2.30 1.95
PCS AUC $\uparrow$ 5 1.20 0.549 0.120 1.480 1.80 2.00 2.20
PROTACs AUC $\uparrow$ 5 2.80 0.247 0.280 1.480 2.60 2.20 1.20
PLI Pearson $\uparrow$ 5 0.40 0.819 0.040 1.480 1.80 2.20 2.00
PFA $F_{\max}\uparrow$ 5 6.40 0.041 0.640 1.480 2.00 2.80 1.20
Notes. $\chi^{2}$ is the Friedman statistic and $W$ is Kendall’s coefficient of concordance ($W=\chi^{2}/(n(k-1))$, with k=3k{=}3 strategies). CD denotes the Nemenyi critical difference at $\alpha{=}0.05$; pairwise average-rank gaps exceeding CD are significant. For PFA, $\lvert\Delta R\rvert=\lvert 1.20-2.80\rvert=1.60>\text{CD}=1.48$ (PFP vs. MVCL). For Overall, the largest rank gap is $\lvert 1.75-2.30\rvert=0.55<0.741$, hence no pairwise difference is significant.
D.3. Additional Observations
To provide a more comprehensive evaluation, we perform full parameter fine-tuning of the pretrained models on three downstream tasks, namely PCS, PROTACs, and PLI, with all encoder parameters updated during training. The corresponding results are presented in the Tab. 6. Generally, performing full-parameter fine-tuning on the entire model with a larger learning rate ($1\times 10^{-4}$) generally leads to slightly lower performance compared to the encoder-freezing setting. This suggests that overly aggressive full fine-tuning is less stable and often degrades ranking and regression metrics across specialized applications. This further reinforces our observation in RQ1: the effective use of pretrained models requires not only preserving pretrained knowledge but also providing the necessary representation adaptation capacity.
Table 10. Results on KIBA Dataset.
EGNN EGNN w/PFP SE(3) Trans SE(3) Trans w/MVCL ProtBERT ProtBERT w/MVCL ESM-2
[…]
@@ D.3. Additional Observations
Figure 7. Visualization of model size and performance across two specialized tasks: (a) PCS and (b) PROTACs. Each point represents a model, with size indicating parameter count and color intensity reflecting task performance. Each point represents a model, with its size corresponding to the number of parameters, ranging from 2M for GVP to 650M for ESM. Larger points indicate larger models. The color intensity of each point indicates performance, with darker shades representing stronger results. The dashed circular boundaries group models by category: models incorporating geometric information in color, protein language models in color, and domain-specific models in color.
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
@@
Figure 8. Visualization of model size and performance across two general tasks: (a) PLI and (b) PFA. Each point represents a model, with size indicating parameter count and color intensity reflecting task performance.
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
Appendix E Preliminaries
E.1. Protein Data
@@
Sequence. The most common representation is the amino acid sequence, typically in FASTA format using single-letter codes. Sequence-based models, originally RNNs and CNNs, are now dominated by Transformers due to their superior long-range modeling. Large pre-trained models such as ESM-2 (Rives et al., 2021; Lin et al., 2023), ProteinBERT (Brandes et al., 2022), and ProtTrans (Elnaggar et al., 2021) capture biochemical properties and evolutionary relationships, with structural information implicitly encoded in their learned representations.
2D Graph.. Proteins can also be described as molecular graphs, with residues as nodes and chemical or spatial relationships as edges. 2D graphs capture connectivity and topological features but lack explicit spatial details. Contact maps and models like GraphFormer (Ying et al., 2021) encode residue-residue proximity or graph connectivity into the attention mechanism, partially reflecting structural information.
3D Structure. Many methods have been developed for modeling the 3D structures of proteins. 3D CNNs employ convolutional neural networks to extract local spatial features of proteins. Graph neural networks are widely used for representing protein and molecular structures, with examples including proteinMPNN (Dauparas et al., 2022), SE(3)-Transformer (Fuchs et al., 2020), GVP (Jing et al., 2021), and GearNet (Zhang et al., 2023). GearNet applies contrastive learning for the pre-training of protein structures. Additionally, Transformer-M (Luo et al., 2023a) incorporates 3D information through distance matrices, enabling transformers to represent protein 3D structures while maintaining invariance. More recently, some works have begun to include protein surface information in modeling; for instance, ProteinINR (Lee et al., 2024) integrates sequence, 3D structure, and surface information to enhance protein representations.
E.2. Terminology
@@
[…]
@@
Appendix F Detail of Downstream Application and Dataset
F.1. Enzyme-Catalyzed Protein Cleavage Site Prediction
Task Definition. Cleavage site prediction is formulated as a large‐scale, imbalanced, residue‐level binary classification problem. Let $P_{s}$ denote a substrate protein of length $\lvert R\rvert$, and for each residue $r_{j}$ ($j=1,\dots,\lvert R\rvert$), let $x_{j}\in\mathbb{R}^{d}$ be its feature vector. The objective is to learn a function
@@ F.1. Enzyme-Catalyzed Protein Cleavage Site Prediction
that outputs $1$ if $r_{j}$ is a cleavage site under catalysis by a given enzyme, and 0 otherwise. Formally, each substrate $P_{s}$ is associated with a label vector $Y\in\{0,1\}^{\lvert R\rvert}$, where $Y_{j}=1$ indicates cleavage at residue $j$. Training proceeds over residue-level examples $\{(x_{j},\,Y_{j})\}$ sampled from annotated cleavage datasets, enabling model generalization to varied substrate contexts.
@@
  • Data Source. The original data are from the MEROPS (Rawlings et al., 2014) database. This dataset has a public website, available at the following address: http://cadd.zju.edu.cn/protacdb/
@@
    • Prior studies (Rawlings et al., 2014) have shown that slight sequence variations among enzymes within the same MEROPS category are negligible. As a result, we generalize the hydrolysis data from a specific substrate-enzyme interaction to all enzymes in that category. This allowed us to augment our dataset by linking each substrate not only to its originally annotated enzyme but also to other enzymes classified in the same MEROPS group. Finally, we selected two enzyme families, C14.003 and M10.003, for evaluation.
@@
    Table 11. Statistics of the dataset after preprocessing.
    Enzyme Family #Train #Test Sites (Train) Sites (Test)
@@
F.2. Targeted Protein Degradation by Proteolysis-Targeting Chimeras
@@
Settings. For this task, we retain the same architectural hyperparameter configuration as used during pretraining. To encode SMILES representations, we adopt the GVP (Jing et al., 2021) architecture with a node dimensionality of 128, an edge dimensionality of 32, and a total of 3 layers. Models are trained for 50 epochs with a learning rate of 5e-4, a batch size of 24, and the Adam optimizer. We report the mean and standard deviation of performance across a fixed set of random seeds. All experiments are conducted on NVIDIA L40 GPUs, with per-run training time ranging from approximately 20 minutes to 6 hours, depending on the model and task complexity.
@@
  • Data Source. The original data was collected from the PROTAC-DB (Ge et al., 2025). This dataset has a public website, available at the following address: http://cadd.zju.edu.cn/protacdb/
@@
    • Following the processing strategy proposed by ET-PROTACs (Cai et al., 2025), we assign a degradation label to each triplet in the dataset. Specifically, we first examine whether a PROTAC entry contains DC50 measurements. If available, entries with DC50 values below 1000 nM are labeled as active, while those equal to or above 1000 nM are labeled as inactive. For entries lacking DC50 data, we instead inspect the reported degradation percentage of the target protein. If available, samples with degradation rates of at least 70% are considered active, and those below 70% are labeled as inactive. In cases where neither DC50 nor degradation percentage is reported, we fall back to IC50 values. Similarly, entries with IC50 below 1000 nM are labeled as active, and those above or equal to this threshold are labeled as inactive.
    • For the target protein and E3 ligase, we retained only the entries with available structures in the AFDB. We filtered out the target proteins with the following UniProt IDs: P36969, P03436, P0DTD1, Q12830, and G0R7E2. The structural data of the target proteins and E3 ligases are obtained from AFDB. Following the approach in (Ingraham et al., 2019), we further convert the format into a dictionary and store it in a .json file. The fields are as follows:
@@
                  "seq":"MADEEAGGTERMEISAELPQTPQRLASWW..",
                  "coord":[[[21.99, 68.31,-49.90], ..],..]
      For each structure, coord is a nested list of shape (N $\times$ 4 $\times$ 3), representing the 3D coordinates of the backbone atoms N, $\alpha, C$, and O for each residue in order, and $N$ is the length of the amino acid sequence.
@@
    uniprot  e3_ligase  linker    warhead     e3_ligand   label
    Q00987    Q96SW2   linker_2  warhead_7   e3_ligand_7   1
    Q00987    Q96SW2   linker_2  warhead_7   e3_ligand_16  1
    P10275    P40337   linker_13 warhead_27  e3_ligand_27  0
    P10275    P40337   linker_33 warhead_27  e3_ligand_27  0
@@
    Table 12. Statistics of PROTACs dataset components.
    Target Protein E3-Ligase Warhead Linker E3-Ligand Label: 1 Label: 0
    154 10 552 1,534 131 2,244 1,878
@@
F.3. Protein–Ligand Interactions